OpenStudy (anonymous):

Cos72=six find the number of degrees in measure of acute angle x 72 is in degree

OpenStudy (dayakar):

\[\cos (90-\theta) =\sin\]

OpenStudy (dayakar):

change 72 into 90-some number

OpenStudy (dayakar):

r u there

OpenStudy (anonymous):

Ya

OpenStudy (dayakar):

can u change

OpenStudy (dayakar):

72 = 90- ?

OpenStudy (dayakar):

can u find the number in the place of ?

OpenStudy (dayakar):

do u understand what i mean

OpenStudy (dayakar):

r u there ,plz respond